ISOLATION AND CHARACTERIZATION OF FORMATE/NI(CYCLAM)^{2+} COMPLEXES WITH CRYO-GENIC ION VIBRATIONAL PREDISSOCIATION

\underline{ARRON B. WOLK}, JOSEPH A. FOURNIER , CONRAD T. WOLKE, and MARK A. JOHNSON, DEPARTMENT OF CHEMISTRY, YALE UNIVERSITY, NEW HAVEN, CT 06520.

Transition metal-based organometallic catalysts are a promising means of converting CO_{2} to transportable fuels. Ni(cyclam)^{2+}(cyclam = 1,4,8,11-tetraazacyclotetradecane), a Ni^{II} complex ligated by four nitrogen centers, has shown promise as a catalyst selective for CO_{2} reduction in aqueous solutions. The cyclam ligand has four NH hydrogen bond donors that can adopt five conformations, each offering distinct binding motifs for coordination of CO_{2} close to the metal center. To probe the ligand conformation and the role of hydrogen bonding in adduct binding, we extract Ni(cyclam)^{2+} complexes with the formate anion and some of its analogs from solution using electrospray ionization, and characterize their structures using cryogenic ion vibrational predissociation spectroscopy. Using the signature vibrational features of the embedded carboxylate anion and the NH groups as reporters, we compare the binding motifs of oxalate, benzoate, and formate anions to the Ni(cyclam)^{2+} framework. Finally, we comment on possible routes to generate the singly charged Ni(cyclam)^{+} complex, a key intermediate that has been invoked in the catalytic CO_{2} reduction cycle, but has never been isolated through ion processing techniques.